OpenStudy (anonymous):

The letters A, B, C, D, E and F are randomly placed in a row. The probability that the letters A and B will occupy the first and second positions respectively is: A 1/15 B 1/3 C 1/30 D 1/6 E 2/3

OpenStudy (ganpat):

letter A occupying 1st seat = 1 out of 6 = 1/6 than letter B occupying second seat = 1 out of 5 = 1/5 so, 1/6 * 1/5 = 1/30.. option C ??

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

how can you differentiate between using permutations nPr or combinatorics nCr?

OpenStudy (ganpat):

Combination is way for selecting things from a group.. for example: selecting a red ball from a group of 15 balls of different colors.. While, Permutation you can say, arrangement of things as per demand.. for example: arrange 10 books, such that 4 & 5 stay together.. This is what i can explain in simplest... Hope u understand..

OpenStudy (ganpat):

http://www.mathsisfun.com/combinatorics/combinations-permutations.html refer this, for understanding...

OpenStudy (anonymous):

okok, thanks.

OpenStudy (ganpat):

anytime :)